A car accelerates uniformly from rest and
reaches a speed of 33.8 m/s in 13.5 s. The
diameter of a tire is 41.5 cm.
Find the number of revolutions the tire
makes during this motion, assuming no slipping.
Answer in units of rev.
Don't round answer
1
Expert's answer
2019-12-17T09:53:41-0500

The distance traveled by car


"d=\frac{v_f+v_i}{2}t=\frac{33.8+0}{2}\times 13.5=228.15\:\rm m"

The number of revolutions the tire makes during this motion


"n=\frac{d}{\pi D}=\frac{228.15}{3.14\times 0.415}=175"